# DNABERT-2: Efficient Foundation Model and Benchmark For Multi-Species Genome

### 26 Jun 2023 v1

@@ Abstract
Decoding the linguistic intricacies of the genome is a crucial problem in biology, and pre-trained foundational models such as DNABERT and Nucleotide Transformer have made significant strides in this area. Existing works have largely hinged on k-mer, fixed-length permutations of A, T, C, and G, as the token of the genome language due to its simplicity. However, we argue that the computation and sample inefficiencies introduced by k-mer tokenization are primary obstacles in developing large genome foundational models. We provide conceptual and empirical insights into genome tokenization, building on which we propose to replace k-mer tokenization with Byte Pair Encoding (BPE), a statistics-based data compression algorithm that constructs tokens by iteratively merging the most frequent co-occurring genome segment in the corpus. We demonstrate that BPE not only overcomes the limitations of k-mer tokenization but also benefits from the computational efficiency of non-overlapping tokenization. Based on these insights, we introduce DNABERT-2, a refined genome foundation model that adapts an efficient tokenizer and employs multiple strategies to overcome input length constraints, reduce time and memory expenditure, and enhance model capability. Furthermore, we identify the absence of a comprehensive and standardized benchmark for genome understanding as another significant impediment to fair comparative analysis. In response, we propose the Genome Understanding Evaluation (GUE), a comprehensive multi-species genome classification dataset that amalgamates $28$ distinct datasets across $7$ tasks, with input lengths ranging from $70$ to $1000$. Through comprehensive experiments on the GUE benchmark, we demonstrate that DNABERT-2 achieves comparable performance to the state-of-the-art model with $21\times$ fewer parameters and approximately $56\times$ less GPU time  [^1] About 14 days on 8 NVIDIA RTX 2080Ti GPUs V.S. 17 days on 128 NVIDIA A100 GPUs. Estimated with the Method 2: GPU Time introduced by OpenAI in https://openai.com/research/ai-and-compute. in pre-training. Compared to DNABERT, while being $3\times$ more efficient, DNABERT-2 outperforms it on $23$ out of $28$ datasets, with an average improvement of $6$ absolute scores on GUE. The code, data, and pre-trained model are publicly available at https://github.com/Zhihan1996/DNABERT_2.
@@ 1 Introduction
Transformer-based foundation models [Bommasani et al., 2022; Kenton and Toutanova, 2019; OpenAI, 2023] have witnessed significant progress in recent years, particularly exemplified by the advent of groundbreaking language models like ChatGPT [Ouyang et al., 2022; OpenAI, 2023]. In parallel, the significance of foundation models has also been increasingly appreciated in the genomics field, as they represent the understanding of genome sequences via numerical embeddings that are directly applicable to various genome analysis tasks. These models can capture complex relationships and dependencies in DNA sequences, opening new avenues for understanding transcriptional regulation [Li et al., 2023], non-coding genetic variants associated with human diseases and traits [Rozowsky et al., 2023], and the functional effects of regulatory elements [Smith et al., 2023]. Recent advancements in genome language modeling have demonstrated their superiority in a range of downstream applications, including promoter prediction [Le et al., 2022; Zhang et al., 2022], gene expression prediction [Avsec et al., 2021], DNA methylation prediction [Jin et al., 2022], chromatin state analysis [Lee et al., 2022], promoter-enhancer interaction prediction [Chen et al., 2022; Ni et al., 2022], TF-DNA binding prediction [Wang et al., 2022], variant effect prediction [Rozowsky et al., 2023], gene network prediction [Theodoris et al., 2023] and more. These models provide researchers with powerful tools to understand the functional importance of different genomics elements and unravel key biological processes and mechanisms.
@@
In this context, we previously developed DNABERT [Ji et al., 2021], an initial foundation model (FM), to unravel the human genome from a language perspective. Despite being widely applied in the community, several technical limitations still present at the time with the original DNABERT implementation, limiting its full potential. First, although proven to be generalizable to other organisms, the pretraining was solely done on the human reference genome, omitting the sequence conservation and diversity across species. Second, k-mer tokenization resulted in information leakage and overall poor computational efficiency during pre-training, which hampers its scalability. Lastly, the simplistic DNABERT-XL solution—intended to bypass the restriction of 512 input sequences imposed by the learned positional embedding [Kenton and Toutanova, 2019]—fell short in handling long input sequences, both in efficiency and effectiveness. These limitations underlined the need for further advancements in the domain of DNA language models.
@@
Recently, Lopez et al. [2023] introduced Nucleotide Transformers (NT), a series of genome foundation models scaling from $500M$ to $2500M$ parameters. NT alleviated the first two limitations of DNABERT by pre-training on a large collection of genomes from 850 species and replacing overlapping k-mer tokenization with a non-overlapping version, substantially reducing tokenized sequence length. Despite this, a hard input length limitation still exist, while, as we will discuss in Sec. 2, non-overlapping k-mer tokenization also suffered from poor sample efficiency as it complicates the model’s task of aligning significantly distinct representations of near-identical inputs.
@@
In view of the aforementioned limitations, we introduce DNABERT-2, a multi-species genome foundation model that replaces k-mer tokenization with Byte Pair Encoding (BPE) [Sennrich et al., 2016], a data compression algorithm that has been widely used by large language models. We show that BPE effectively addresses the known issues of k-mer tokenization while maintaining the computational efficiency of non-overlapping tokenization. Moreover, DNABERT-2 overcomes the limitation of DNABERT by replacing learned positional embeddings with Attention with Linear Biases (ALiBi) [Press et al., 2021] to get rid of the input length limitation, incorporating Flash Attention [Dao et al., 2022] to increase computational efficiency, and adjusting model architecture to increase model capability. As a result of the efficient tokenizer and advanced model architecture, DNABERT-2 achieves comparable performance to the state-of-the-art model with approximately $56\times$ less computational cost and $21\times$ fewer parameters, identifying its computation- and sample- efficiency and enabling efficient fine-tuning on most consumer GPUs.
@@
Meanwhile, despite progress in genome foundational models, the absence of carefully curated benchmarks has posed a significant challenge. Owing to the unstandardized pre-processing pipeline of genome sequences, it is unjust to directly compare model performances with results reported in previous papers, even when the data originate from the same source. Moreover, many genome understanding evaluation datasets used in existing works [Lopez et al., 2023] are either too trivial or too challenging, leading to similar scores for most models and failing to accurately reflect different models’ capabilities. The scarcity of high-quality benchmark datasets hampers evaluating and comparing different models and further hinders the development of novel techniques. To this end, we introduce Genome Understanding Evaluation (GUE), a standardized and comprehensive multi-species benchmark containing $28$ datasets across $7$ important genome analysis tasks on genomes of $4$ species with input lengths ranging from $70$ to $1000$. All the datasets are elaborately calibrated with a series of strategies to ensure they are suitable for reflecting the capability level of existing genome foundation models.
@@
Our main contributions can be therefore summarized as follows: 1) We identify key obstacles in genome tokenization and provide deep insights, presenting a simple yet effective solution that balances the efficiency and effectiveness of genome foundation models; 2) We introduce DNABERT-2, an efficient pre-trained foundation model for multi-species genome that delivers performance on par with the state-of-the-art model while being $21\times$ smaller and utilizes approximately $56\times$ less GPU time; 3) We introduce Genome Understanding Evaluation (GUE), a standardized, comprehensive, and well-calibrated multi-species genome classification benchmark including $7$ tasks and $28$ datasets to facilitate research in genome foundation model.
@@ 2 Background
Tokenization serves as a critical initial step in language modeling, significantly impacting the efficiency and effectiveness of the model. DNA sequences consist of $4$ unique nucleotide bases: A, T, C, and G. A majority of genome language models [Ji et al., 2021; Lopez et al., 2023] utilize the k-mer tokenization technique, in which each contiguous $k$-length genome segment is considered as a token. During tokenization, a sliding window with window size $k$ and stride $t$ is employed to convert the original genome sequence into a series of k-mers. Here, the stride $t$ is either set as $1$ or $k$, while the first one represents the overlapping version of k-mer tokenization and the other one represents the non-overlapping version. Figure 1 presents examples of overlapping (left) and non-overlapping (right) k-mer tokenizations. Despite its wide application, we argue that both versions of the k-mer tokenization are less optimal.
@@
Overlapping k-mers tokenization ensures adjacent tokens always overlap by k1k-1 characters, resulting in significant information leakage in masked language modeling. As depicted in Figure 1, a masked token is entirely leaked when adjacent tokens from both sides are not masked, and it is partially leaked when adjacent tokens from only one side are present. Generally, in the overlapping $k$-mer tokenization setting, let $l$ and $r$ denote the distances between a masked token [M] and its closest unmasked adjacent token on the left and right sides, the number of possible options of [M] is $4^{\min(l,r,k,\max(0,l+r-k))}$. In other words, to prevent the entire leakage of a masked token, at least k1k-1 tokens on its left and right sides in total must be masked, which explains why Ji et al. [2021] opt to mask a continuous span of $k$ tokens. Furthermore, to guarantee no leakage of a masked token, at least $k$ tokens on both sides must be masked. Nevertheless, information leakage is still inevitable for the leftmost and rightmost k1k-1 masked tokens. Ideally, in masked language modeling, a model is required to select the best option from the entire vocabulary, enabling it to differentiate and evaluate among a large number of options. However, if the search space is undesirably reduced due to information leakage, the model only needs to differentiate between a limited number of options. Consequently, this results in poor sample efficiency, as the model may not be sufficiently challenged to learn the underlying patterns in the data. Also, the tokenized sequence for an input of length $L$ consists of Lk+1L-k+1 tokens, each with a length of $k$. This results in a tokenized sequence with considerable redundancy and a length nearly equivalent to the original sequence, leading to low computation efficiency considering the quadratic computation complexity of Transformer-based [Vaswani et al., 2017] models. This becomes particularly problematic when attempting to scale up the model. Therefore, Lopez et al. [2023] proposed the non-overlapping k-mer tokenization.
@@
To address the aforementioned issues, we propose to adapt SentencePiece [Kudo and Richardson, 2018], a subword tokenization framework widely used in natural language processing, to replace k-mer tokenization for genome sequences. We employ Byte-Pair Encoding (BPE) [Sennrich et al., 2016] to iteratively merge frequent pairs of nucleotides and genome segments, forming a vocabulary of variable-length tokens that effectively represent the entire genome dataset. Despite its conceptual simplicity, this method is well-suited for genome foundation models. First, it not only prevents information leakage but also significantly reduces the sequence length by approximately $5$ times (detailed statistics are presented in Sec 3.1), substantially improving computational efficiency. Moreover, its robust tokenization result is beneficial for sample efficiency since it allows the model to focus on understanding the genome language semantics without being distracted by the distinct representations of the same input. Furthermore, unlike k-mer tokenization, BPE doesn’t always produce tokens of length $k$. Consequently, when a token containing an unspecified number of nucleotides is masked, the model is challenged to predict both the number of nucleotides and the particular nucleotides themselves. This naturally transforms the masked language modeling objective into a T5-style [Raffel et al., 2020] "replace spans of text" objective, which has been demonstrated to be more effective than standard masked language modeling in various scenarios.
@@ 3 Method
In this section, we provide empirical analysis on the BPE tokenizer for genome language (§ 3.1) and describe the model architecture (§ 3.2) and implementation details (§ 3.3) of DNABERT-2.
@@ 3.1 Tokenizer
DNABERT-2 adapts SentencePiece [Kudo and Richardson, 2018] with Byte Pair Encoding (BPE) [Sennrich et al., 2016] to perform tokenization for DNA sequences. SentencePiece is a language-agnostic tokenizer that considers each input as a raw stream without assuming any pre-tokenization, which matches greatly with genome sequences where the definitions of word and sentence do not exist. BPE is a compression algorithm that has been widely used in the area of natural language processing as a word segmentation strategy. It learns a fixed-sized vocabulary of variable-length tokens based on the co-occurrence frequency of the characters. Figure 2 illustrates the process of constructing a vocabulary from a given corpus with BPE. First, we initialize the vocabulary with all unique characters in the corpus. Then, in each iteration, we view the most frequent character segment (e.g., TA at iteration $1$) as a new word, add it to the vocabulary, and update the corpus by replacing all the same segments with this new word. The iteration continues till we achieve the desired number of words in the vocabulary. Thus, the target vocabulary size plays a crucial role.
Due to the significant difference between natural language and DNA sequence, vocabulary sizes that are commonly used in the NLP area [Kenton and Toutanova, 2019; Vaswani et al., 2017; Raffel et al., 2020; OpenAI, 2023] may not be appropriate for genome sequences. To determine the most suitable vocabulary size, we constructed $8$ vocabularies with target sizes ranging from $2^{8}$ to $2^{15}$ on the multi-species genomes (see Sec. 4.1) to empirically evaluate the impact of varying vocabulary sizes. As indicated in Figure 3a, larger vocabularies tend to encompass more lengthy tokens, which enables the tokenizer to represent the same input sequence with fewer tokens. Shorter tokenized sequences consequently reduce the computational cost (See Figure 3b), as the computational complexity of Transformers is quadratic in relation to the input sequence length. Therefore, from the computation efficiency perspective, a larger vocabulary size is favorable.

### 18 Mar 2024 v2

@@ Abstract
Decoding the linguistic intricacies of the genome is a crucial problem in biology, and pre-trained foundational models such as DNABERT and Nucleotide Transformer have made significant strides in this area. Existing works have largely hinged on k-mer, fixed-length permutations of A, T, C, and G, as the token of the genome language due to its simplicity. However, we argue that the computation and sample inefficiencies introduced by k-mer tokenization are primary obstacles in developing large genome foundational models. We provide conceptual and empirical insights into genome tokenization, building on which we propose to replace k-mer tokenization with Byte Pair Encoding (BPE), a statistics-based data compression algorithm that constructs tokens by iteratively merging the most frequent co-occurring genome segment in the corpus. We demonstrate that BPE not only overcomes the limitations of k-mer tokenization but also benefits from the computational efficiency of non-overlapping tokenization. Based on these insights, we introduce DNABERT-2, a refined genome foundation model that adapts an efficient tokenizer and employs multiple strategies to overcome input length constraints, reduce time and memory expenditure, and enhance model capability. Furthermore, we identify the absence of a comprehensive and standardized benchmark for genome understanding as another significant impediment to fair comparative analysis. In response, we propose the Genome Understanding Evaluation, a comprehensive multi-species genome classification dataset that amalgamates $36$ distinct datasets across $9$ tasks, with input lengths ranging from $70$ to $10000$. Through comprehensive experiments on the GUE benchmark, we demonstrate that DNABERT-2 achieves comparable performance to the state-of-the-art model with $21\times$ fewer parameters and approximately $92\times$ less GPU time  [^1] About 14 days on 8 NVIDIA RTX 2080Ti V.S. 28 days on 128 NVIDIA A100. Estimated with the Method 2: GPU Time introduced by OpenAI in https://openai.com/research/ai-and-compute. in pre-training. Compared to DNABERT, while being $3\times$ more efficient, DNABERT-2 outperforms it on $23$ out of $28$ datasets, with an average improvement of $6$ absolute scores on GUE. The code, data, and pre-trained model are available at https://github.com/MAGICS-LAB/DNABERT_2.
@@ 1 Introduction
Transformer-based foundation models (Bommasani et al., 2022; Kenton & Toutanova, 2019; OpenAI, 2023) have witnessed significant progress in recent years, particularly exemplified by the advent of groundbreaking language models like ChatGPT (Ouyang et al., 2022; OpenAI, 2023). In parallel, the significance of foundation models has also been increasingly appreciated in the genomics field, as they represent the understanding of genome sequences via numerical embeddings that are directly applicable to various genome analysis tasks. These models can capture complex relationships and dependencies in DNA sequences, opening new avenues for understanding transcriptional regulation (Li et al., 2023), non-coding genetic variants associated with human diseases and traits (Rozowsky et al., 2023), and the functional effects of regulatory elements (Smith et al., 2023). Recent advancements in genome language modeling have demonstrated their superiority in a range of downstream applications, including promoter prediction (Le et al., 2022; Zhang et al., 2022), gene expression prediction (Avsec et al., 2021), DNA methylation prediction (Jin et al., 2022), chromatin state analysis (Lee et al., 2022), promoter-enhancer interaction prediction (Chen et al., 2022; Ni et al., 2022), TF-DNA binding prediction (Wang et al., 2022), variant effect prediction (Rozowsky et al., 2023), gene network prediction (Theodoris et al., 2023) and more. These models provide researchers with powerful tools to understand the functional importance of different genomics elements and unravel key biological processes and mechanisms.
@@
In this context, Ji et al. (2021) developed DNABERT, an initial foundation model (FM), to unravel the human genome from a language perspective. Despite being widely applied in the community, several technical limitations still present at the time with the original DNABERT implementation, limiting its full potential. First, although proven to be generalizable to other organisms, the pretraining was solely done on the human reference genome, omitting the sequence conservation and diversity across species. Second, k-mer tokenization resulted in information leakage and overall poor computational efficiency during pre-training, which hampers its scalability. Lastly, the simplistic DNABERT-XL solution—intended to bypass the restriction of 512 input sequences imposed by the learned positional embedding (Kenton & Toutanova, 2019)—fell short in handling long input sequences, both in efficiency and effectiveness. These limitations underlined the need for further advancements in the domain of DNA language models.
@@
Recently, Dalla-Torre et al. (2023) introduced Nucleotide Transformers (NT), a series of genome foundation models scaling from $500M$ to $2500M$ parameters. NT alleviated the first two limitations of DNABERT by pre-training on a large collection of genomes from 850 species and replacing overlapping k-mer tokenization with a non-overlapping version, substantially reducing tokenized sequence length. Despite this, a hard input length limitation still exist, while, as we will discuss in Sec. 2, non-overlapping k-mer tokenization also suffered from poor sample efficiency as it complicates the model’s task of aligning significantly distinct representations of near-identical inputs.
@@
In view of the aforementioned limitations, we introduce DNABERT-2, a multi-species genome foundation model that replaces k-mer tokenization with Byte Pair Encoding (BPE) (Sennrich et al., 2016), a data compression algorithm that has been widely used by large language models. We show that BPE effectively addresses the known issues of k-mer tokenization while maintaining the computational efficiency of non-overlapping tokenization. Moreover, DNABERT-2 overcomes the limitation of DNABERT by replacing learned positional embeddings with Attention with Linear Biases (ALiBi) (Press et al., 2021) to get rid of the input length limitation, incorporating Flash Attention (Dao et al., 2022) to increase computational efficiency, and adjusting model architecture to increase model capability. As a result of the efficient tokenizer and advanced model architecture, DNABERT-2 achieves comparable performance to the state-of-the-art model with approximately $92\times$ less computational cost and $21\times$ fewer parameters, identifying its computation- and sample- efficiency and enabling efficient fine-tuning on most consumer GPUs.
@@
Meanwhile, despite progress in genome foundational models, the absence of carefully curated benchmarks has posed a significant challenge. Owing to the unstandardized pre-processing pipeline of genome sequences, it is unjust to directly compare model performances with results reported in previous papers, even when the data originate from the same source. Moreover, many genome understanding evaluation datasets used in existing works (Dalla-Torre et al., 2023) are either too trivial or too challenging, leading to similar scores for most models and failing to accurately reflect different models’ capabilities. The scarcity of high-quality benchmark datasets hampers evaluating and comparing different models and further hinders the development of novel techniques. To this end, we introduce Genome Understanding Evaluation (GUE & GUE+), a standardized and comprehensive multi-species benchmark containing $36$ datasets across $9$ important genome analysis tasks on genomes of $4$ species with input lengths ranging from $70$ to $10000$. All the datasets are elaborately calibrated with a series of strategies to ensure they are suitable for reflecting the capability level of existing genome foundation models.
@@
Our main contributions can be therefore summarized as follows: 1) We identify key obstacles in genome tokenization and provide deep insights, presenting a simple yet effective solution that balances the efficiency and effectiveness of genome foundation models; 2) We introduce DNABERT-2, an efficient pre-trained foundation model for multi-species genome that delivers performance on par with the state-of-the-art model while being $21\times$ smaller and utilizes approximately $92\times$ less GPU time; 3) We introduce Genome Understanding Evaluation (GUE), a standardized, comprehensive, and well-calibrated multi-species genome classification benchmark including $9$ tasks and $36$ datasets to facilitate research in genome foundation model.
@@ 2 Background
Tokenization serves as a critical initial step in language modeling, significantly impacting the efficiency and effectiveness of the model. DNA sequences consist of $4$ unique nucleotide bases: A, T, C, and G. A majority of genome language models (Ji et al., 2021; Dalla-Torre et al., 2023) utilize the k-mer tokenization technique, in which each contiguous $k$-length genome segment is considered as a token. During tokenization, a sliding window with window size $k$ and stride $t$ is employed to convert the original genome sequence into a series of k-mers. Here, the stride $t$ is either set as $1$ or $k$, while the first one represents the overlapping version of k-mer tokenization and the other one represents the non-overlapping version. Figure 1 presents examples of overlapping (left) and non-overlapping (right) k-mer tokenizations. Despite its wide application, we argue that both versions of the k-mer tokenization are less optimal.
@@
Overlapping k-mers tokenization ensures adjacent tokens always overlap by k1k-1 characters, resulting in significant information leakage in masked language modeling. As depicted in Figure 1, a masked token is entirely leaked when adjacent tokens from both sides are not masked, and it is partially leaked when adjacent tokens from only one side are present. Generally, in the overlapping $k$-mer tokenization setting, let $l$ and $r$ denote the distances between a masked token [M] and its closest unmasked adjacent token on the left and right sides, the number of possible options of [M] is $4^{\min(l,r,k,\max(0,l+r-k))}$. In other words, to prevent the entire leakage of a masked token, at least k1k-1 tokens on its left and right sides in total must be masked, which explains why Ji et al. (2021) opt to mask a continuous span of $k$ tokens. Furthermore, to guarantee no leakage of a masked token, at least $k$ tokens on both sides must be masked. Nevertheless, information leakage is still inevitable for the leftmost and rightmost k1k-1 masked tokens. Ideally, in masked language modeling, a model is required to select the best option from the entire vocabulary, enabling it to differentiate and evaluate among a large number of options. However, if the search space is undesirably reduced due to information leakage, the model only needs to differentiate between a limited number of options. Consequently, this results in poor sample efficiency, as the model may not be sufficiently challenged to learn the underlying patterns in the data. Also, the tokenized sequence for an input of length $L$ consists of Lk+1L-k+1 tokens, each with a length of $k$. This results in a tokenized sequence with considerable redundancy and a length nearly equivalent to the original sequence, leading to low computation efficiency considering the quadratic computation complexity of Transformer-based (Vaswani et al., 2017) models. This becomes particularly problematic when attempting to scale up the model. Therefore, Dalla-Torre et al. (2023) proposed the non-overlapping k-mer tokenization.
@@
To address the aforementioned issues, we propose to adapt SentencePiece (Kudo & Richardson, 2018), a subword tokenization framework widely used in natural language processing, to replace k-mer tokenization for genome sequences. We employ Byte-Pair Encoding (BPE) (Sennrich et al., 2016) to iteratively merge frequent pairs of nucleotides and genome segments, forming a vocabulary of variable-length tokens that effectively represent the entire genome dataset. Despite its conceptual simplicity, this method is well-suited for genome foundation models. First, it not only prevents information leakage but also significantly reduces the sequence length by approximately $5$ times (detailed statistics are presented in Sec 3.1), substantially improving computational efficiency. Moreover, its robust tokenization result is beneficial for sample efficiency since it allows the model to focus on understanding the genome language semantics without being distracted by the distinct representations of the same input. Furthermore, unlike k-mer tokenization, BPE doesn’t always produce tokens of length $k$. Consequently, when a token containing an unspecified number of nucleotides is masked, the model is challenged to predict both the number of nucleotides and the particular nucleotides themselves. This naturally transforms the masked language modeling objective into a T5-style (Raffel et al., 2020) "replace spans of text" objective, which has been demonstrated to be more effective than standard masked language modeling in various scenarios.
@@ 3 Method
In this section, we provide empirical analysis on the BPE tokenizer for genome language (§ 3.1) and describe the model architecture (§ 3.2) and implementation details (§ A.2) of DNABERT-2.
@@ 3.1 Tokenizer
DNABERT-2 adapts SentencePiece (Kudo & Richardson, 2018) with Byte Pair Encoding (BPE) (Sennrich et al., 2016) to perform tokenization for DNA sequences. SentencePiece is a language-agnostic tokenizer that considers each input as a raw stream without assuming any pre-tokenization, which matches greatly with genome sequences where the definitions of word and sentence do not exist. BPE is a compression algorithm that has been widely used in the area of natural language processing as a word segmentation strategy. It learns a fixed-sized vocabulary of variable-length tokens based on the co-occurrence frequency of the characters. Figure 2 illustrates the process of constructing a vocabulary from a given corpus with BPE. First, we initialize the vocabulary with all unique characters in the corpus. Then, in each iteration, we view the most frequent character segment (e.g., TA at iteration $1$) as a new word, add it to the vocabulary, and update the corpus by replacing all the same segments with this new word. The iteration continues till we achieve the desired number of words in the vocabulary. Thus, the target vocabulary size plays a crucial role.
Due to the significant difference between natural language and DNA sequence, vocabulary sizes that are commonly used in the NLP area (Kenton & Toutanova, 2019; Vaswani et al., 2017; Raffel et al., 2020; OpenAI, 2023) may not be appropriate for genome sequences. To determine the most suitable vocabulary size, we constructed $8$ vocabularies with target sizes ranging from $2^{8}$ to $2^{15}$ on the multi-species genomes (see Sec. 4.1) to empirically evaluate the impact of varying vocabulary sizes. As indicated in Figure 3a, larger vocabularies tend to encompass more lengthy tokens, which enables the tokenizer to represent the same input sequence with fewer tokens. Shorter tokenized sequences consequently reduce the computational cost (See Figure 3b), as the computational complexity of Transformers is quadratic in relation to the input sequence length. Therefore, from the computation efficiency perspective, a larger vocabulary size is favorable.